OpenStudy (anonymous):

Factor Completely: 3x2 - 21 *3(x2 - 7)?

OpenStudy (ash2326):

Here we have \[3x^2-21\] \[3(x^2-7)\] We can write \[7= ( \sqrt 7)^2\] so we have \[3(x^2- (\sqrt 7)^2)\] Can you solve now? Use \[a^2-b^2=(a+b)(a-b)\]

OpenStudy (ash2326):

@Haleybob Did you try?